# L(M)V-IQL: Multiple Intention Inverse Reinforcement Learning for Animal Behavior Characterization

### v1

@@ A.1 Proof of Theorem 1
Substitute the parameters $\Theta=\{\nu_{1},\ldots,\nu_{K};{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$ under independent latent state assumption into the EM update equation (Equation 4) and unroll:
@@
$=\sum_{{\mathcal{Y}}}\sum_{i=1}^{N}\log(\nu_{y_{i}}\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{y_{i}}}))\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$ (A.1)
$=\sum_{y_{1}}\cdots\sum_{y_{N}}\sum_{i=1}^{N}\sum_{k=1}^{K}\delta_{k=y_{i}}\log(\nu_{k}\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$ (A.2)
$=\sum_{k=1}^{K}\sum_{i=1}^{N}\log(\nu_{k}\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\sum_{y_{1}}\cdots\sum_{y_{N}}\delta_{k=y_{i}}\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$ (A.3)
$=\sum_{k=1}^{K}\sum_{i=1}^{N}\log(\nu_{k}\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\zeta_{ik}^{\tau}$ (by Equation 5)
$=\sum_{k=1}^{K}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\log(\nu_{k})+\sum_{k=1}^{K}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\text{\,,}$ (A.4)
where $\delta$ denotes the Kronecker delta function. Equation A.4 indicates that $\nu_{k}$ and ${\mathcal{R}}_{k}$ are not interdependent, we can thus optimize them separately in the M-step of EM, leading to the second update equation in Equation 5 trivially. According to Gibbs’ inequality, the first term of Equation A.4 is maximized if and only if
@@ A.2 Proof of Theorem 2
Similar to the proof for Theorem 1, substitute the parameter set $\Theta=\{\Pi,\Lambda;{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\}$ into the EM update equation (Equation 4) and unroll:
@@
$=\sum_{{\mathcal{Y}}}\log(\pi_{y_{0}}\Pr(\xi_{0}\mid\pi_{{\mathcal{R}}_{y_{0}}})\prod_{i=1}^{N}\Lambda_{y_{i-1}y_{i}}\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{y_{i}}}))\Pr({\mathcal{Y}}\mid{\mathcal{D}},\Theta^{\tau})$ (A.6)
@@
$\quad+\sum_{{\mathcal{Y}}}\sum_{i=0}^{N}\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{y_{i}}}))\Pr({\mathcal{Y}}\mid{\mathcal{D}},\Theta^{\tau})$ (A.7)
@@
$\quad+\sum_{y_{1}}\cdots\sum_{y_{N}}\sum_{i=0}^{N}\sum_{k=1}^{K}\delta_{k=y_{i}}\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$ (A.8)
@@
$\quad+\sum_{k=1}^{K}\sum_{i=0}^{N}\Pr(y_{i}=k\mid{\mathcal{D}},\Theta^{\tau})\log(\Pr(\xi_{i}\mid\pi_{{\mathcal{R}}_{k}}))\text{\,,}$ (A.9)
where $\delta$ denotes the Kronecker delta function. Since $\pi_{k}$, $\Lambda_{kl}$ and ${\mathcal{R}}_{k}$ are not interdependent, we can thus maximize the respective term separately, resulting in Equation 6. ∎
Remark 1.
In a more practical case where the agent demonstration space has multiple trajectory sequences, Theorem 2 can also be generalized and proved in the same manner.
Remark 2.
All $\xi\in{\mathcal{D}}$ above are assumed to be the trajectory for a whole episode. In some special cases where it is assumed that the latent state transition happens after each action execution, instead of per episode, Theorem 2 can also be applied by regarding each episode as a trajectory sequence with each trajectory consists of only one action execution.
@@ B.1 Latent Variable Inverse Q-learning
1 initialize $\Theta\coloneqq\{\nu_{1},\ldots,\nu_{K};{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$ repeat
2       E-step
3             foreach $\xi_{i}\in{\mathcal{D}}$ do
4                   forall $k$ do
5                         $\zeta_{ik}\leftarrow\prod_{(s,a)\in\xi_{i}}\pi_{{\mathcal{R}}_{k}}(s,a)\nu_{k}/Z$
6
7
8
9      M-step
10             forall $k$ do
11                   $\nu_{k}\leftarrow\sum_{i=1}^{N}\zeta_{ik}/N$ compute ${\mathcal{R}}_{k}$ via IAVI or IQL on ${\mathcal{D}}$ with weight $\zeta_{ik}$ on trajectory $\xi_{i}$
12
13
14until convergence;
@@ B.2 Latent Markov Variable Inverse Q-learning
$&\Pi_{k}^{\tau+1}\coloneqq\mathfrak{g}_{0k}^{\tau}\\ &\Lambda_{kl}^{\tau+1}\coloneqq\frac{\sum_{i=1}^{N}\mathfrak{x}_{ikl}^{\tau}}{\sum_{i=0}^{N-1}\mathfrak{g}_{ik}^{\tau}}\\ &{\mathcal{R}}_{k}^{\tau+1}\coloneqq\operatorname*{arg\,max}_{{\mathcal{R}}_{k}}\sum_{i=0}^{N}\mathfrak{g}_{ik}^{\tau}\log(\Pr(\xi_{i}\mid\Pi_{{\mathcal{R}}_{k}}))\text{\,.}\end{aligned}\right.$ (B.5)
@@
Input: expert demonstrations ${\mathcal{D}}$, latent space dimension $K$
1 initialize $\Theta\coloneqq\{\Pi,\Lambda;{\mathcal{R}}_{1},\ldots,{\mathcal{R}}_{K}\}$ repeat
2       E-step
3             calculate $\mathfrak{g}$ and $\mathfrak{x}$ according to Equations B.1B.4
4
5      M-step
6             forall $k$ do
7                   $\Pi_{k}\leftarrow\mathfrak{g}_{0k}$ forall $l$ do
8                         $\Lambda_{kl}\leftarrow\sum_{i=1}^{N}\mathfrak{x}_{ikl}/\sum_{i=0}^{N-1}\mathfrak{g}_{ik}$
9                  compute ${\mathcal{R}}_{k}$ via IAVI or IQL on ${\mathcal{D}}$ with weight $\mathfrak{g}_{ik}$ on trajectory $\xi_{i}$
10
11
12until convergence;
@@ 8                         $\Lambda_{kl}\leftarrow\sum_{i=1}^{N}\mathfrak{x}_{ikl}/\sum_{i=0}^{N-1}\mathfrak{g}_{ik}$
Algorithm 2 Latent Markov Variable Inverse Q-learning (LMV-IQL)
@@ C.2 Additional Results on Gridworld
We also performed analysis under the environment set up where the intention irrelevant punishments were removed, i.e. replacing the $-1$ reward on the type of reward irrelevant to the intentions with 00. In this environment, there is an increased overlapping between some of the demonstrated trajectories under different intentions (Figure 6). However, LV-IAVI and LV-IQL still outperform the single intention algorithm IAVI and IQL in trajectory clustering and recovering corresponding expert reward functions.
@@ D.2 Additional Results for Water-unrestricted Mice
Figure 7: (A) Comparison of LMV-IAVI on test set trajectories to a random policy and DIRL, represented as LL. (B) Learnt policy (red arrows) under different intentions and corresponding state occupancy (grey colormap) in the environment. State occupancy was calculated by assigning each trajectory to the latent state with hightest posterior probability. Policies are shown only for some states. (C) Trajectories of latent state probabilities. olid and shaded curves denote the Mean and SE. (D) Inferred latent state transition matrix from the best-fitting LMV-IAVI.
In contrast to the outcomes from the water-restricted animal dataset (Figure 3), LMV-IAVI demonstrates a higher test LL, even when considering a single intention. As the number of latent states (associated with DIRL’s goal maps) grows, the test LL of LMV-IAVI increases, while the test LL of DIRL remains constant (Figure 7). Focus on the two latent states LMV-IAVI, the inferred policy under two intentions exhibits ‘Exploring’ and ‘Tired’ behavior. The policy under ‘Exploring’ tends to encourage the animal lingering in the labyrinth, whereas the policy under ‘Tired’ latent state steers the animal back to the maze entrance. Correspondingly, the posterior probability of ‘Exploring’ initially dominates at the session’s beginning but is generally surpassed by the ‘Tired’ latent state over time.
@@ Appendix E Further Details on the Evaluation of Mice Reversal-learning Behavior
1 foreach $\ell_{h}\in\{1,\ldots,5\}$ do
2       run IQL on the concatenated data from all animals until convergence
select best $\ell_{h}$ with largest test set LL Fit LMV-IQL globally:
3 foreach $K\in\{2,\ldots,5\}$ do
4       foreach $i\in\{1,\ldots,10\}$ do
5             initialize LMV-IQL with $K$ latent states and random parameters run Algorithm 2 on the concatenated data from all animals until convergence
6
@@
7 forall animals do
8       foreach $K\in\{1,\ldots,5\}$ do
9             initialize LMV-IQL with $K$ latent states using the best global fit parameters for this $K$ run Algorithm 2 until convergence
10
Algorithm 3 Fitting LMV-IQL on real mice behavior
fBwB4ngcAcByHJpPJl+fn54mD3Gg0NrquXxeLRQAAwzAYj8cwTZPwPH9/sVg8PXweDAauqqr2cDjEer1GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@ Appendix A Proof of Theorems
Substitute the parameters $\Theta=\{\nu_{1},\ldots,\nu_{K};r_{1},\ldots,r_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$ under independent latent state assumption into the EM update equation (Equation 5) and unroll:
@@
$=\sum_{\mathcal{Y}}\sum_{i=1}^{N}\log(\nu_{y_{i}}\Pr(\xi_{i}\mid\pi_{r_{y_{i}}}))\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$ (A.1)
$=\sum_{y_{1}}\cdots\sum_{y_{N}}\sum_{i=1}^{N}\sum_{k=1}^{K}\delta_{k=y_{i}}\log(\nu_{k}\Pr(\xi_{i}\mid\pi_{r_{k}}))\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$ (A.2)
$=\sum_{k=1}^{K}\sum_{i=1}^{N}\log(\nu_{k}\Pr(\xi_{i}\mid\pi_{r_{k}}))\sum_{y_{1}}\cdots\sum_{y_{N}}\delta_{k=y_{i}}\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$ (A.3)
$=\sum_{k=1}^{K}\sum_{i=1}^{N}\log(\nu_{k}\Pr(\xi_{i}\mid\pi_{r_{k}}))\zeta_{ik}^{\tau}$ (by Equation 7)
$=\sum_{k=1}^{K}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\log(\nu_{k})+\sum_{k=1}^{K}\sum_{i=1}^{N}\zeta_{ik}^{\tau}\log(\Pr(\xi_{i}\mid\pi_{r_{k}}))\text{\,,}$ (A.4)
where $\delta$ denotes the Kronecker delta function. Equation A.4 indicates that $\nu_{k}$ and $r_{k}$ are not interdependent, we can thus optimize them separately in the M-step of EM, leading to the second update equation in Equation 6 trivially. According to Gibbs’ inequality, the first term of Equation A.4 is maximized if and only if
@@
Similar to the proof for Theorem 4.2, substitute the parameter set $\Theta=\{\Pi,\Lambda;r_{1},\ldots,r_{K}\}$ into the EM update equation (Equation 5) and unroll:
@@
$=\sum_{\mathcal{Y}}\log(\Pi_{y_{0}}\Pr(\xi_{0}\mid\pi_{r_{y_{0}}})\prod_{i=1}^{N}\Lambda_{y_{i-1}y_{i}}\Pr(\xi_{i}\mid\pi_{r_{y_{i}}}))\Pr(\mathcal{Y}\mid\mathcal{D},\Theta^{\tau})$ (A.6)
@@
$\quad+\sum_{\mathcal{Y}}\sum_{i=0}^{N}\log(\Pr(\xi_{i}\mid\pi_{r_{y_{i}}}))\Pr(\mathcal{Y}\mid\mathcal{D},\Theta^{\tau})$ (A.7)
@@
$\quad+\sum_{y_{1}}\cdots\sum_{y_{N}}\sum_{i=0}^{N}\sum_{k=1}^{K}\delta_{k=y_{i}}\log(\Pr(\xi_{i}\mid\pi_{r_{k}}))\prod_{i^{\prime}=1}^{N}\Pr(y_{i^{\prime}}\mid\xi_{i^{\prime}},\Theta^{\tau})$ (A.8)
@@
$\quad+\sum_{k=1}^{K}\sum_{i=0}^{N}\Pr(y_{i}=k\mid\mathcal{D},\Theta^{\tau})\log(\Pr(\xi_{i}\mid\pi_{r_{k}}))\text{\,,}$ (A.9)
where $\delta$ denotes the Kronecker delta function. Since $\Pi_{k}$, $\Lambda_{kl}$ and $r_{k}$ are not interdependent, we can thus maximize the respective term separately, resulting in Equation 8. ∎
Remark A.1.
In a more practical case where the agent demonstration space has multiple trajectory sequences, Theorem 4.3 can also be generalized and proved in the same manner.
Remark A.2.
All $\xi\in\mathcal{D}$ above are assumed to be the trajectory for a whole episode. In some special cases where it is assumed that the latent state transition happens after each action execution, instead of per episode, Theorem 4.3 can also be applied by regarding each episode as a trajectory sequence with each trajectory consists of only one action execution.
@@ B.1 Latent Variable Inverse Q-learning
  Initialize $\Theta\coloneqq\{\nu_{1},\ldots,\nu_{K};r_{1},\ldots,r_{K}\mid\nu_{1}+\cdots+\nu_{K}=1\}$.
  repeat
     E-step:
     for each $\xi_{i}\in\mathcal{D}$ do
        for all $k$ do
           $\zeta_{ik}\leftarrow\prod_{(s,a)\in\xi_{i}}\pi_{r_{k}}(s,a)\nu_{k}/Z$
        end for
     end for
     M-step:
     for all $k$ do
        $\nu_{k}\leftarrow\sum_{i=1}^{N}\zeta_{ik}/N$
        Compute $r_{k}$ via IAVI or IQL on $\mathcal{D}$ with weight $\zeta_{ik}$ on trajectory $\xi_{i}$.
     end for
  until convergence
@@ B.2 Latent Markov Variable Inverse Q-learning
$&\Pi_{k}^{\tau+1}\coloneqq\mathfrak{g}_{0k}^{\tau}\\ &\Lambda_{kl}^{\tau+1}\coloneqq\frac{\sum_{i=1}^{N}\mathfrak{x}_{ikl}^{\tau}}{\sum_{i=0}^{N-1}\mathfrak{g}_{ik}^{\tau}}\\ &r_{k}^{\tau+1}\coloneqq\operatorname*{argmax}_{r_{k}}\sum_{i=0}^{N}\mathfrak{g}_{ik}^{\tau}\log(\Pr(\xi_{i}\mid\pi_{r_{k}}))\text{\,.}\end{aligned}\right.$ (B.5)
@@
Algorithm 2 Latent Markov Variable Inverse Q-learning (LMV-IQL)
  Input: agent demonstrations $\mathcal{D}$, latent space dimension $K$
  Initialize $\Theta\coloneqq\{\Pi,\Lambda;r_{1},\ldots,r_{K}\}$.
  repeat
     E-step:
     Calculate $\mathfrak{g}$ and $\mathfrak{x}$ according to Equations B.1 to B.4.
     M-step:
     for all $k$ do
        $\Pi_{k}\leftarrow\mathfrak{g}_{0k}$
        for all $l$ do
           $\Lambda_{kl}\leftarrow\sum_{i=1}^{N}\mathfrak{x}_{ikl}/\sum_{i=0}^{N-1}\mathfrak{g}_{ik}$
        end for
        Compute $r_{k}$ via IAVI or IQL on $\mathcal{D}$ with weight $\mathfrak{g}_{ik}$ on trajectory $\xi_{i}$.
     end for
  until convergence
@@ C.2 Additional Results on Gridworld
Figure C.1: (Top) Visualization of the normalized ground truth and learnt state-value functions. Red lines indicate the ground truth trajectory distribution and the learnt trajectory clusters used to recover the reward function for respective intention. (All expert trajectories are shown for each figure of IAVI and IQL.) (Bottom) EVD for different approaches, Mean $\pm$ SE over 5 repeated runs.
We also performed analysis under the environment set up where the intention irrelevant punishments were removed, i.e. replacing the $-1$ reward on the type of reward irrelevant to the intentions with 00. In this environment, there is an increased overlapping between some of the demonstrated trajectories under different intentions (Figure C.1). However, LV-IAVI and LV-IQL still outperform the single intention algorithm IAVI and IQL in trajectory clustering and recovering corresponding expert reward functions.
@@ D.2 Additional Results for Water-unrestricted Mice
Figure D.1: (A) Comparison of LMV-IAVI on test set trajectories to a random policy and DIRL, represented as LL. (B) Learnt policy (red arrows) under different intentions and corresponding state occupancy (grey colormap) in the environment. State occupancy was calculated by assigning each trajectory to the latent state with hightest posterior probability. Policies are shown only for some states. (C) Trajectories of latent state probabilities. olid and shaded curves denote the Mean and SE. (D) Inferred latent state transition matrix from the best-fitting LMV-IAVI.
In contrast to the outcomes from the water-restricted animal dataset (Figure 3), LMV-IAVI demonstrates a higher test LL, even when considering a single intention. As the number of latent states (associated with DIRL’s goal maps) grows, the test LL of LMV-IAVI increases, while the test LL of DIRL remains constant (Figure D.1). Focus on the two latent states LMV-IAVI, the inferred policy under two intentions exhibits ‘Exploring’ and ‘Tired’ behavior. The policy under ‘Exploring’ tends to encourage the animal lingering in the labyrinth, whereas the policy under ‘Tired’ latent state steers the animal back to the maze entrance. Correspondingly, the posterior probability of ‘Exploring’ initially dominates at the session’s beginning but is generally surpassed by the ‘Tired’ latent state over time.
@@ Appendix E Further Details on the Evaluation of Mice Reversal-learning Behavior
  for each $\ell_{h}\in\{1,\ldots,5\}$ do
     Run IQL on the concatenated data from all animals until convergence.
  end for
  Select best $\ell_{h}$ with largest test set log-likelihood.
  Fit LMV-IQL globally:
  for each $K\in\{2,\ldots,5\}$ do
     for each $i\in\{1,\ldots,10\}$ do
        Initialize LMV-IQL with $K$ latent states and random parameters.
        Run Algorithm 2 on the concatenated data from all animals until convergence.
     end for
  end for
@@
  for all animals do
     for each $K\in\{1,\ldots,5\}$ do
        Initialize LMV-IQL with $K$ latent states using the best global fit parameters for this $K$.
        Run Algorithm 2 until convergence.
     end for
  end for